OpenStudy (anonymous):

find the cartesian equation of r = 8 sin thet + 8 cos theta

OpenStudy (amistre64):

its a circle centered at the irigin of radius 8 maybe?

OpenStudy (anonymous):

\[r =\sqrt{x ^{2}+y ^{2}}\]

OpenStudy (amistre64):

polars and parametrics aint my strong point ;)

OpenStudy (anonymous):

youre good at vectors :)

OpenStudy (amistre64):

yep, I can point at things all day lol

OpenStudy (anonymous):

haha, and good at planes

OpenStudy (anonymous):

maybe youre good at visualizing things?

OpenStudy (amistre64):

im pretty good at visualizing stuff

OpenStudy (anonymous):

ok we can use x = r cos theta, and y = r sin theta,

OpenStudy (anonymous):

so multiply both sides by r

OpenStudy (amistre64):

spinning polars tho.....not so much; aint had the practice

OpenStudy (anonymous):

\[\sin \theta =x/r\]

OpenStudy (anonymous):

ive read the vector stuff, it just wont stick. for some reason

OpenStudy (amistre64):

polars are just vector equations at heart :)

OpenStudy (anonymous):

i dont see that

OpenStudy (amistre64):

r = magnitude; <cos,sin> are the components

OpenStudy (amistre64):

r<cos(t),sin(t)> is the basi set up

OpenStudy (anonymous):

those are the cartesian components you mean

OpenStudy (amistre64):

in the plane, yes

OpenStudy (amistre64):

but polars define length(r) whih is the magnitude of a vector; and the <cos,sin> angles are the x and y components of a vector

OpenStudy (amistre64):

a vector function simply defines a curve or surface generated by the parametric equations for the vector components from teh origin

OpenStudy (amistre64):

and that is all a polar equation is

OpenStudy (anonymous):

come again, parametric equation for the vector components (the cartesian components?)

OpenStudy (amistre64):

(r,t) is a polar equation right? (radius,theta) this tells you how far to turn and how far to move

OpenStudy (amistre64):

thats all a vector is; an arrow indicating direction and length

OpenStudy (anonymous):

ok , lets use th for theta

OpenStudy (anonymous):

ok , so say again your statement

OpenStudy (amistre64):

which one lol

OpenStudy (amistre64):

r<cos(th),sin(th)> is the vector equivalent of a polar equation (r,th)

OpenStudy (amistre64):

or simpy <r cos(th), r sin(th)>

OpenStudy (anonymous):

vector , as in the cartesian components of the vector

OpenStudy (amistre64):

yes

OpenStudy (amistre64):

the point P(x,y) is the same as defining a vector from the origin as <x,y>

OpenStudy (amistre64):

the vector is an arrow pointing to the point

OpenStudy (anonymous):

right, that sometimes confuses me

OpenStudy (anonymous):

we write < x,y> for a vector, and P(x,y) for a point

OpenStudy (amistre64):

sometimes they write a vector as v(x,y) which confuses tha tmatter; i prefer the convention of just making it pointy to indicate its an arrow :)

OpenStudy (anonymous):

right, i like to distinguish between points (n tuples) and vectors

myininaya (myininaya):

cantorset i posted a proof for your viewing sorry it took me awhile to respond

OpenStudy (anonymous):

i cant find it, one sec

myininaya (myininaya):

k